OpenStudy (anonymous):

solve for x 3[2-4(2x-1)]=4x-10

OpenStudy (anonymous):

3[2-4(2x-1)]=4x-10 6-12(2x-1)=4x-10 6-24x+12=4x-10

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

yyes it does thank you... i messed up by also distributing the 3 to the -1

OpenStudy (precal):

 alt way to do problem always work from the inside out ie distribute the 4 first (please)